# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ 6.2 Voxelwise Encoding Model
The main goal of the voxel-wise encoding model is to predict brain responses associated with each brain voxel given a stimulus. To estimate the brain alignment of a DNN model of stimulus representations via training standard voxel-wise encoding models Deniz et al. (2019); Toneva & Wehbe (2019). Specifically, for each voxel and participant, prior studies train fMRI encoding model using ridge regression to predict the fMRI recording associated with this voxel as a function of the stimulus representations obtained from DNN models. To simultaneously accommodate different feature spaces, which may necessitate varying levels of regularization, Nunez-Elizalde et al. (2019) proposed voxel-wise encoding model that utilize an advanced form of ridge regression. This method, known as banded ridge regression, introduces individual regularization parameters for each feature space, as illustrated in Figure 11. Before doing the ridge regression or banded ridge regression, we first z-scored each feature channel separately for training and testing. This was done to match the features to the fMRI responses, which were also z-scored for training and testing. Formally, at the time step (t), we encode the stimuli as $X_{t}\in\mathbb{R}^{N\times D}$ and brain region voxels $Y_{t}\in\mathbb{R}^{N\times V}$, where $N$ is the number of training examples, $D$ denotes the dimension of the concatenation of delayed TRs, and $V$ denotes the number of voxels. To find the optimal regularization parameter for each feature space, we use a range of regularization parameters that is explored using cross-validation.
@@ 6.3 Interpreting brain recordings through the robustness of DNN model representations
Variance partitioning. Variance partitioning quantifies the unique contribution of different stimulus features to BOLD responses. For variance partitioning (Lescroart, 2017; Deniz et al., 2019; Vaidya et al., 2022), set theory is used to calculate the common variance (as the intersection of various combinations of feature spaces) and the unique variance (as the set difference for each individual feature space), as shown in Figure 12 (a). Overall, the total variance explained by each model is computed as the unique variance explained by each model and the shared variance across models. This variance partition approach was computed separately for each voxel, then averaged across ROIs and across subjects.
Indirect approach. An indirect approach first relates model representations to the human brain, followed by an independent examination of the related model to some task performance or behavioral output. For instance, Schrimpf et al. (2021) tests the computations of a language model that may underlie human language understanding. This is accomplished by an independent examination of the relationship between the models’ ability to predict an upcoming word and their brain predictivity. Similarly, Goldstein et al. (2022) provides empirical evidence that both the human brain and language model engage in continuous next-word prediction before word onset.
Residual approach. In contrast to indirect approach, the approach proposed in Toneva et al. (2022a) can directly estimate the impact of a specific feature on the alignment between the model and the brain recordings by observing the difference in alignment before and after the specific feature is computationally removed from the model representations. This method use to remove the linear contribution of a feature to a model’s representation is one way to implement such a direct approach, as shown in Figure 12 (b). This is why residual approach also refer to as direct. Another method was investigated by previous work Oota et al. (2024b); Dong & Toneva (2023) and was shown to yield very similar results.
@@
Stacked regression. The stacked regression approach, proposed by Lin et al. (2023), follows a two-level pipeline. The first level consists of different linear regressors, each using a different stimulus feature space as input. At the second level, the parameters $\alpha_{j}$ are learned for a convex combination of first level predictors. Overall, the entire stacked model is estimated separately at each voxel. This method is useful when building different encoding models where input feature spaces are correlated (e.g., visual and semantic features of natural images) and for demonstrating the importance of each feature space in predicting a voxel’s response.
@@
Caucheteux et al. (2021) GPT-2 345 Narratives 7.5secs (5 TRs)
@@
Stimuli Authors Dataset Type Lang. Stimulus Representations ||||S|||| Dataset
@@
(Toneva & Wehbe, 2019) fMRI / MEG English ELMo, BERT, Transformer-XL 9 Story understanding
(Toneva et al., 2020) MEG English BERT 9 Question-Answering
(Schrimpf et al., 2021) fMRI / ECoG English 43 language models (e.g. GloVe, ELMo, BERT, GPT-2, XLNET) 20 Neural architecture of language
(Gauthier & Levy, 2019) fMRI English BERT, finetuned NLP tasks (Sentiment, Natural language inference), Scrambling language model 7 Imagine a situation with the noun
@@
 (Caucheteux et al., 2021) fMRI English GPT-2, Basic syntax features 345 Narratives
@@
(Goldstein et al., 2022) fMRI English GloVe, GPT-2 next word, pre-onset, post-onset word surprise 8 ECoG
@@
 (Oota et al., 2022a) fMRI English ESN, LSTM, ELMo, Longformer 82 Narratives
(Merlin & Toneva, 2022) fMRI English BERT, Next word prediction, multi-word semantics, scrambling model 8 Harry Potter
(Toneva et al., 2022a) fMRI / MEG English ELMo, BERT, Context Residuals 8 Harry Potter
(Aw & Toneva, 2023) fMRI English BART, Longformer, Long-T5, BigBird, and corresponding Booksum models as well 8 Passive reading
(Zhang et al., 2022b) fMRI English, Chinese Node Count 19, 12 Zhang
@@
(Oota et al., 2023b) MEG English Basic syntax features, GloVe and BERT 8 MEG-MASC
(Tuckute et al., 2024) fMRI English BERT-Large, GPT-2 XL 12 Reading Sentences
@@
(Singh et al., 2023) fMRI English BERT-Large, GPT-2 XL, Text Perturbations 5 Pereira
@@
(Kubilius et al., 2019) fMRI - CNN models AlexNet, ResNet, DenseNet 7 Algonauts
@@
(Conwell et al., 2023) fMRI - CNN models AlexNet 4 BOLD 5000
@@
(Oota et al., 2023d) fMRI English Wav2Vec2.0 and SUPERB tasks 82 Narratives
@@
(Oota et al., 2022e) fMRI English CLIP, VisualBERT, LXMERT, CNNs and BERT 5, 82 Periera & Narratives
@@ 6.4 Linguistic Encoding
6.4.1 Alignment Between Pretrained Language Models (LMs) and Brains
@@
Figure 13: (a) Alignment of representations between deep learning systems and human brains (Toneva & Wehbe, 2019). (b) For instance, a narrative story provided to both the Language model as well as human participants. For the Language model, we extract its representations for every word in the text. For the human participants, we record their brain activity using fMRI. Next, we train a linear function that uses the extracted Language model representations to predict human brain activity. Finally, we test this function on unseen data, and evaluate its accuracy as the amount of “brain alignment” (Toneva & Wehbe, 2019). These two images are sourced from Cogsci-22 tutorial slides Oota et al. (2022d).
@@
Previous works have investigated the alignment between pretrained language models and brain recordings of people comprehending language.  Huth et al. (2016) have been able to identify brain ROIs (Regions of Interest) that respond to words that have a similar meaning and have thus built a “semantic atlas” of how the human brain organizes language. Many studies have shown accurate results in mapping brain activity using neural distributed word embeddings for linguistic stimuli  (Anderson et al., 2017a; Pereira et al., 2018; Oota et al., 2018; Nishida & Nishimoto, 2018; Sun et al., 2019). Unlike earlier models, where each word is represented as an independent vector in an embedding space, Jain & Huth (2018) built encoding models using rich contextual representations derived from an LSTM language model in a story listening task. With these contextual representations, demonstrated dissociation in brain activation – auditory cortex (AC) and Broca’s area in shorter context whereas left Temporo-Parietal junction (TPJ) in longer context.  Hollenstein et al. (2019) presents the first multimodal framework for evaluating six types of word embeddings (Word2Vec, WordNet2Vec (Bartusiak et al., 2019), GloVe, fastText, ELMo, and BERT) on 15 datasets, including eye-tracking, EEG and fMRI signals recorded during language processing. With the recent advances in contextual representations in NLP, few studies incorporated them in relating sentence embeddings with brain activity patterns (Sun et al., 2020; Gauthier & Levy, 2019; Jat et al., 2020).
@@
Figure 14: The strongest alignment with high-level language brain regions has consistently been observed in the middle layers. Left: Performance of BERT encoding model for all hidden layers as the amount of context provided to the network is increased (Toneva & Wehbe, 2019). Right: fMRI encoding score (averaged across time and channels) of 6 representative transformers varying in tasks (CLM vs MLM) and depth (4-12 layers) (Caucheteux & King, 2020).  The left Figure is adapted from Toneva & Wehbe (2019) and the right Figure is adapted from Caucheteux & King (2020).
@@
More recently, researchers have begun to study the alignment of language regions of the brain with the layers of language models (broadly following the method described in Figure 13) and found that the best alignment was achieved in the middle layers of these models (Jain & Huth, 2018; Toneva & Wehbe, 2019; Caucheteux & King, 2020), as shown in Figure 14.  Toneva & Wehbe (2019) study how representations of various Transformer models differ across layer depth, context length, and attention type. The results demonstrated that across several larger NLP models, middle layers of language models are well aligned with brain language regions. Schrimpf et al. (2021) examined the relationship between 43 diverse state-of-the-art language models. They also studied the behavioral signatures of human language processing in the form of self-paced reading times and a range of linguistic functions assessed via standard engineering tasks from NLP. They found that Transformer-based models perform better than RNNs or word-level embedding models. Larger-capacity models perform better than smaller models. Models initialized with random weights (prior to training) perform surprisingly similarly in neural predictivity compared to final trained models, suggesting that network architecture contributes as much or more than experience dependent learning to a model’s match to the brain. Antonello et al. (2021) proposed a “language representation embedding space” and demonstrated the effectiveness of the features from this embedding in predicting fMRI responses to linguistic stimuli. Very recent work by (Antonello et al., 2024) tested whether larger open-source models, such as those from the text-based model (OPT and LLaMA) families, are better at predicting brain responses recorded using fMRI. The results demonstrate that encoding performance improvements scale well with both model size and dataset size, and large datasets will no doubt be necessary in producing useful encoding models.
6.4.2 Disentangling the Syntax and Semantics
@@
The representations of transformer models like BERT and GPT-2 have been shown to linearly map onto brain activity during language comprehension. Several studies have attempted to disentangle the contributions of different types of information from word representations to the alignment between brain recordings and language models (Lopopolo et al., 2017; Wang et al., 2020; Caucheteux et al., 2021; Reddy & Wehbe, 2021; Zhang et al., 2022a; Toneva et al., 2022a; Oota et al., 2023a).  Wang et al. (2020) proposed a two-channel variational autoencoder model to dissociate sentences into semantic and syntactic representations and separately associate them with brain imaging data to find feature-correlated brain regions. Similarly,  Zhang et al. (2022a) separated different syntactic features from pretrained BERT representations, to explore the potential for distinct syntactic and semantic processing language regions in the brain. Compared to lexical word representations, word syntactic features (parts-of-speech, named entities) and word-relation features (semantic roles, dependencies) are distributed across brain networks instead of a local brain region. The previous two studies could not conclude whether all or any of these representations effectively drive the linear mapping between language models (LMs) and the brain.  Toneva et al. (2022a) presented an approach to disentangle supra-word meaning from lexical meaning in language models and showed that supra-word meaning is predictive of fMRI recordings in two language regions (anterior and posterior temporal lobes). Similar to the approach presented in Toneva et al. (2022a)Oota et al. (2023b) disentangle the past and future context meaning from word meaning in language models and showed that past context is crucial in obtaining significant results while predicting MEG brain recordings.  Caucheteux et al. (2021) proposed a taxonomy to factorize the high-dimensional activations of language models into four combinatorial classes: lexical, compositional, syntactic, and semantic representations. They found that (1) Compositional representations recruit a more widespread cortical network than lexical ones and encompass the bilateral temporal, parietal, and prefrontal cortices. (2) Contrary to previous claims, syntax and semantics are not associated with separated modules, but, instead, appear to share a common and distributed neural substrate.
@@
Figure 15: Four steps proposed in (Oota et al., 2023a): (1) fMRI acquisition, (2) Syntactic parsing, (3) Regression model training, and (4) Predictive power analysis of the three embeddings methods. This Figure is adapted from Oota et al. (2023a).
@@
While previous works studied syntactic processing as captured through complexity measures (syntactic surprisal, node count, word length, and word frequency) (Zhang et al., 2020; 2022a), very few have studied the syntactic representations themselves (Caucheteux et al., 2021; Reddy & Wehbe, 2021; Oota et al., 2023a). Studying syntactic representations using fMRI is difficult because (1) representing syntactic structure in an embedding space is a non-trivial computational problem, and (2) the fMRI signal is noisy. To overcome these limitations, Reddy & Wehbe (2021) proposed syntactic structure embeddings that encode the syntactic information inherent in the natural text that subjects read in the scanner. The results reveal that syntactic structure-based features explain additional variance in the brain activity of various parts of the language system, even after controlling for complexity metrics that capture the processing load. Toneva et al. (2022b) further examined whether the representations obtained from a language model align with different language processing regions in a similar or different way. While Reddy & Wehbe (2021) focused on constituency parsing mainly including incremental top-down parsing, Oota et al. (2023a) leverage dependency information more systematically by learning the dependency representations using graph convolutional networks, using the four step recipe as illustrated in Figure 15. The results reveal that constituency tree structure is better encoded in language regions such as bilateral temporal cortex (ATL and PTL) and MFG, while dependency structure is better encoded in AG and PCC language regions.
@@
While previous studies focused on narrative English language stories and have shown that several brain regions are involved in building the hierarchical syntactic structure, a recent study in (Zhang et al., 2022b) analyzes the neural basis of such structures between two diverse languages: Chinese and English. The results demonstrate that the brain may use different parsing strategies for different language structures to reduce the cognitive load.

### v3

@@ 6.2 Voxelwise Encoding Model
The main goal of the voxel-wise encoding model is to predict brain responses associated with each brain voxel given a stimulus. To estimate the brain alignment of a DNN model of stimulus representations, researchers train standard voxel-wise encoding models (Deniz et al., 2019; Toneva & Wehbe, 2019). Specifically, for each voxel and participant, prior studies train fMRI encoding models using ridge regression to predict the fMRI recording associated with this voxel as a function of the stimulus representations obtained from DNN models. To simultaneously accommodate different feature spaces, which may necessitate varying levels of regularization, Nunez-Elizalde et al. (2019) proposed a voxel-wise encoding model that utilizes an advanced form of ridge regression. This method, known as banded ridge regression, introduces individual regularization parameters for each feature space, as illustrated in Figure 10 (b). Before performing ridge regression or banded ridge regression, we first z-score normalize each feature channel separately for training and testing. This is done to match the features to the fMRI responses, which are also z-scored for training and testing. Formally, at time step (t), we encode the stimuli as $X_{t}\in\mathbb{R}^{N\times D}$ and brain region voxels $Y_{t}\in\mathbb{R}^{N\times V}$, where $N$ is the number of training examples, $D$ denotes the dimension of the concatenation of delayed TRs, and $V$ denotes the number of voxels. To find the optimal regularization parameter for each feature space, we use a range of regularization parameters that is explored using cross-validation.
@@ 6.3 Interpreting brain recordings through the robustness of DNN model representations
Variance partitioning. Variance partitioning quantifies the unique contribution of different stimulus features to BOLD responses. For variance partitioning (Lescroart, 2017; Deniz et al., 2019; Vaidya et al., 2022), set theory is used to calculate the common variance (as the intersection of various combinations of feature spaces) and the unique variance (as the set difference for each individual feature space), as shown in Figure 11 (a). Overall, the total variance explained by each model is computed as the unique variance explained by each model and the shared variance across models. This variance partition approach is computed separately for each voxel, then averaged across ROIs and across subjects.
Task-Performance aligned approach. The task-Performance aligned approach first relates model representations to the human brain data, followed by an independent examination of the related model to some task performance or behavioral output. For instance, Schrimpf et al. (2021) tests the computations of a language model that may underlie human language understanding. This is accomplished by an independent examination of the relationship between the models’ ability to predict an upcoming word and their brain predictivity. Similarly,Goldstein et al. (2022) provides empirical evidence that both the human brain and language models engage in continuous next-word prediction before word onset. Analogous to linguistic encoding studies, many vision encoding studies adopt this approach to validate the significance of relating task performance to encoding performance (Yamins & DiCarlo, 2016; Schrimpf et al., 2020; Cao et al., 2021).
Residual approach. In contrast to the indirect approach, the method proposed by Toneva et al. (2022a) can directly estimate the impact of a specific feature on the alignment between the model and the brain recordings by observing the difference in alignment before and after the specific feature is computationally removed from the model representations. This method, used to remove the linear contribution of a feature to a model’s representation, is one way to implement such a direct approach, as shown in Figure 11 (b). This is why the residual approach is also referred to as direct. Another method was investigated by previous work Oota et al. (2024b); Dong & Toneva (2023) and was shown to yield very similar results.
@@
Wehbe et al. (2014) Non-negative sparse embedding 8 Harry-Potter 8secs (4 TRs)
@@
Caucheteux et al. (2021a) GPT-2 345 Narratives 7.5secs (5 TRs)
@@
Chen et al. (2024a) BERT 9 Moth-Radio-Hour 8secs (4 TRs)
Chen et al. (2024b) mBERT 6 Bilingual Moth-Radio-Hour 8secs (4 TRs)
@@
Stacked regression. The stacked regression approach, proposed by Lin et al. (2023), follows a two-level pipeline. The first level consists of different linear regressors, each using a different stimulus feature space as input. At the second level, the parameters $\alpha_{j}$ are learned for a convex combination of first-level predictors. Overall, the entire stacked model is estimated separately at each voxel. This method is useful when building different encoding models where input feature spaces are correlated (e.g., visual and semantic features of natural images) and for demonstrating the importance of each feature space in predicting a voxel’s response.
Hierarchical processing in DNNs and the Human Brain. Hierarchical processing is a fundamental characteristic observed in both biological neural systems and deep neural networks (DNNs). To investigate the hierarchy of information processing, researchers primarily focus on an in-depth analysis of how hierarchical structures in deep neural networks (DNNs) mirror the hierarchical processing observed in the human visual cortex, as well as in speech and language comprehension. For example, Yamins & DiCarlo (2016) demonstrate that early layers in convolutional neural networks (CNNs) correspond to primary visual areas (e.g., V1, V2) that process basic features, while deeper layers align with higher visual areas (e.g., IT cortex) responsible for complex object recognition. Similarly, Caucheteux et al. (2021b) employ a model-based approach to successfully replicate the seminal study by Lerner et al. (2011), which revealed the hierarchy of language areas by comparing fMRI data of subjects listening to both regular and scrambled narratives.
Authors Dataset Type Lang. Stimulus Representations ||||S|||| Dataset
@@
(Toneva & Wehbe, 2019) fMRI, MEG English ELMo, BERT, Transformer-XL 9 Harry Potter
(Jat et al., 2020) MEG English ELMo, BERT 8 Rafidi2016 MEG data
(Schrimpf et al., 2021) fMRI, ECoG English 43 language models (e.g. GloVe, ELMo, BERT, GPT-2, XLNET) 20 Blank2014, Fedorenko2016, Pereira2018
(Gauthier & Levy, 2019) fMRI English BERT, finetuned NLP tasks (Sentiment, Natural language inference), Scrambling language model 8 Pereira2018
@@
(Zhang et al., 2020) fMRI English Word2Vec 19 Moth Radio Hour
 (Caucheteux et al., 2021a) fMRI English GPT-2, Basic syntax features 345 Narratives
@@
(Goldstein et al., 2022) fMRI English GloVe, GPT-2 next word, pre-onset, post-onset word surprise 8 Goldstein2022 ECoG
@@
 (Lamarre et al., 2022) fMRI English BERT, Attention Heads 9 Subset Moth Radio Hour
(Toneva et al., 2022a) fMRI, MEG English ELMo, BERT, Context Residuals 8 Harry Potter
(Aw & Toneva, 2023) fMRI English BART, Longformer, Long-T5, BigBird, and corresponding Booksum models as well 8 Harry Potter
(Zhang et al., 2022c) fMRI English, Chinese Node Count 19, 12 Zhang
@@
(Chen et al., 2024a) fMRI English Time scale features of BERT 9 Subset Moth Radio Hour
(Tuckute et al., 2024) fMRI English BERT-Large, GPT-2 XL 12 Tuckute2024
@@
(Singh et al., 2023) fMRI English BERT-Large, GPT-2 XL, Text Perturbations 5 Pereira2018
@@
(Kubilius et al., 2019) fMRI - CNN models AlexNet, ResNet, DenseNet 1472 Human behavioral data
@@
(Conwell et al., 2024) fMRI - CNN models AlexNet 4 BOLD 5000
@@
(Li et al., 2024b) fMRI English Wav2Vec2.0 and SUPERB tasks 21 Narratives
@@
(Wang et al., 2023) fMRI English CLIP 5 BOLD5000
@@ 6.4 Linguistic Encoding
6.4.1 Alignment Between Pretrained Language Models (LMs) and the Brain
@@
Figure 12: (a) Alignment of representations between deep learning systems and human brains (Toneva & Wehbe, 2019). (b) For instance, a narrative story provided to both the Language model as well as human participants. For the Language model, we extract its representations for every word in the text. For the human participants, we record their brain activity using fMRI. Next, we train a linear function that uses the extracted Language model representations to predict human brain activity. Finally, we test this function on unseen data, and evaluate its accuracy as the amount of “brain alignment” (Toneva & Wehbe, 2019). Here, LM refers to the Language Model, and SM refers to the Speech Model. These two images are sourced from Cogsci-22 tutorial slides Oota et al. (2022d), and used with permission from the respective authors.
@@
Previous works have investigated the alignment between pretrained language models and brain recordings of people comprehending language. Huth et al. (2016) were able to identify brain ROIs (Regions of Interest) that respond to words with similar meanings, thus building a “semantic atlas” of how the human brain organizes language. Many studies have shown accurate results in mapping brain activity using neural distributed word embeddings for linguistic stimuli  (Anderson et al., 2017a; Pereira et al., 2018; Oota et al., 2018; Nishida & Nishimoto, 2018; Sun et al., 2019). Unlike earlier models, where each word is represented as an independent vector in an embedding space, Jain & Huth (2018) built encoding models using rich contextual representations derived from an LSTM language model in a story listening task. Using these contextual representations, their approach revealed a clear dissociation dissociation in brain activation, with the auditory cortex (AC) and Broca’s area processing shorter contexts, whereas the left Temporo-Parietal junction (TPJ) became more engaged with longer contexts.  Hollenstein et al. (2019) presents the first multimodal framework for evaluating six types of word embeddings (Word2Vec, WordNet2Vec (Bartusiak et al., 2019), GloVe, fastText, ELMo, and BERT) on 15 datasets, including eye-tracking, EEG and fMRI signals recorded during language processing. With the recent advances in contextual representations in NLP, few studies incorporated them in relating sentence embeddings with brain activity patterns (Sun et al., 2020; Gauthier & Levy, 2019; Jat et al., 2020).
@@
Figure 13: The strongest alignment with high-level language brain regions has consistently been observed in the middle layers. Left: Performance of BERT encoding model for all hidden layers as the amount of context provided to the network is increased (Toneva & Wehbe, 2019). Right: fMRI encoding score (averaged across time and channels) of 6 representative transformers varying in tasks (CLM vs MLM) and depth (4-12 layers) (Caucheteux & King, 2022).  The left Figure is adapted from (Toneva & Wehbe, 2019) and the right Figure is adapted from (Caucheteux & King, 2022). Adapted with permission from the respective authors. Note: Please see a cautionary note about the robustness of middle-layer alignment in the main text.
@@
More recently, researchers have begun to study the alignment of language regions of the brain with the layers of language models (broadly following the method described in Figure 12) and found that the best alignment was achieved in the middle layers of these models (Jain & Huth, 2018; Toneva & Wehbe, 2019; Caucheteux & King, 2022), as shown in Figure 13.  Toneva & Wehbe (2019) study how representations of various Transformer models differ across layer depth, context length, and attention type. The results demonstrated that across several larger NLP models, middle layers of language models are well aligned with brain language regions. However, this higher brain alignment in the middle layers is influenced by the model’s architecture type (encoder, decoder, or encoder-decoder) rather than being a universal characteristic. Schrimpf et al. (2021) examined the relationship between 43 diverse state-of-the-art language models. They also studied the behavioral signatures of human language processing in the form of self-paced reading times and a range of linguistic functions assessed via standard engineering tasks from NLP. They found that Transformer-based models perform better than RNNs or word-level embedding models. Larger-capacity models perform better than smaller models. Models initialized with random weights (prior to training) perform surprisingly similarly in neural predictivity compared to final trained models, suggesting that network architecture contributes as much or more than experience dependent learning to a model’s match to the brain. Antonello et al. (2021) proposed a “language representation embedding space” and demonstrated the effectiveness of the features from this embedding in predicting fMRI responses to linguistic stimuli. Very recent work by Antonello et al. (2024b) tested whether larger open-source models, such as those from the text-based model (OPT (Zhang et al., 2022a) and LLaMA (Touvron et al., 2023)) families, are better at predicting brain responses recorded using fMRI. The results demonstrate that encoding performance improvements scale well with both model size and dataset size, and large datasets will no doubt be necessary in producing useful encoding models.
6.4.2 Disentangling Syntax and Semantics
@@
The representations of transformer models like BERT and GPT-2 have been shown to linearly map onto brain activity during language comprehension. Several studies have attempted to disentangle the contributions of different types of information from word representations to the alignment between brain recordings and language models (Lopopolo et al., 2017; Wang et al., 2020; Caucheteux et al., 2021a; Reddy & Wehbe, 2021; Zhang et al., 2022b; Toneva et al., 2022a; Oota et al., 2023a).  Wang et al. (2020) proposed a two-channel variational autoencoder model to dissociate sentences into semantic and syntactic representations and separately associate them with brain imaging data to find feature-correlated brain regions. Similarly,  Zhang et al. (2022b) separated different syntactic features from pretrained BERT representations, to explore the potential for distinct syntactic and semantic processing language regions in the brain. Compared to lexical word representations, word syntactic features (parts-of-speech, named entities) and word-relation features (semantic roles, dependencies) are distributed across brain networks instead of a local brain region. The previous two studies could not conclude whether all or any of these representations effectively drive the linear mapping between language models (LMs) and the brain.  Toneva et al. (2022a) presented an approach to disentangle supra-word meaning from lexical meaning in language models and showed that supra-word meaning is predictive of fMRI recordings in two language regions (anterior and posterior temporal lobes). Similar to the approach presented in Toneva et al. (2022a)Oota et al. (2023b) disentangle the past and future context meaning from word meaning in language models and showed that past context is crucial in obtaining significant results while predicting MEG brain recordings.  Caucheteux et al. (2021a) proposed a taxonomy to factorize the high-dimensional activations of language models into four combinatorial classes: lexical, compositional, syntactic, and semantic representations. They found that (1) Compositional representations recruit a more widespread cortical network than lexical ones and encompass the bilateral temporal, parietal, and prefrontal cortices. (2) Contrary to previous claims, syntax and semantics are not associated with separated modules, but, instead, appear to share a common and distributed neural substrate.
@@
Figure 14: Four steps proposed in Oota et al. (2023a): (1) fMRI acquisition, (2) Syntactic parsing, (3) Regression model training, and (4) Predictive power analysis of the three embedding methods. This Figure is adapted from (Oota et al., 2023a), and used with permission from the respective authors.
@@
While previous works studied syntactic processing as captured through complexity measures (syntactic surprisal, node count, word length, and word frequency) (Zhang et al., 2020; 2022b), very few have studied the syntactic representations themselves (Caucheteux et al., 2021a; Reddy & Wehbe, 2021; Oota et al., 2023a). Studying syntactic representations using fMRI is difficult because (1) representing syntactic structure in an embedding space is a non-trivial computational problem, and (2) the fMRI signal is noisy. To overcome these limitations, Reddy & Wehbe (2021) proposed syntactic structure embeddings that encode the syntactic information inherent in the natural text that subjects read in the scanner. The results reveal that syntactic structure-based features explain additional variance in the brain activity of various parts of the language system, even after controlling for complexity metrics that capture the processing load. Toneva et al. (2022b) further examined whether the representations obtained from a language model align with different language processing regions in a similar or different way. While Reddy & Wehbe (2021) focused on constituency parsing mainly including incremental top-down parsing, Oota et al. (2023a) leverage dependency information more systematically by learning the dependency representations using graph convolutional networks, using the four step recipe as illustrated in Figure 14. The results reveal that constituency tree structure is better encoded in language regions such as bilateral temporal cortex (ATL and PTL) and MFG, while dependency structure is better encoded in AG and PCC language regions.
@@
While previous studies focused on narrative English language stories and have shown that several brain regions are involved in building the hierarchical syntactic structure, a recent study by Zhang et al. (2022c) analyzes the neural basis of such structures between two diverse languages: Chinese and English. The results demonstrate that the brain may use different parsing strategies for different language structures to reduce the cognitive load.